### v2

# SwitchHead: Accelerating Transformers with Mixture-of-Experts Attention

Abstract
The costly self-attention layers in modern Transformers require memory and compute quadratic in sequence length. Existing approximation methods usually underperform and fail to obtain significant speedups in practice. Here we present SwitchHead—a novel method that reduces both compute and memory requirements and achieves wall-clock speedup, while matching the language modeling performance of baseline Transformers with the same parameter budget. SwitchHead uses Mixture-of-Experts (MoE) layers for the value and output projections and requires 4 to 8 times fewer attention matrices than standard Transformers. Our novel attention can also be combined with MoE MLP layers, resulting in an efficient fully-MoE “SwitchAll” Transformer model. Our code is public. [^1] https://github.com/robertcsordas/moe_attention ${}^{\dagger}$ork done at IDSIA.
1 Introduction
Large language models (LLMs) have shown remarkable capabilities (Radford et al., 2019; Brown et al., 2020; OpenAI, 2022; 2023) and great versatility (Bubeck et al., 2023). However, training enormous Transformers (Vaswani et al., 2017; Schmidhuber, 1992) requires a considerable amount of computing power and memory, which is not accessible to most researchers, academic institutions, and even companies. Even running them in inference mode, which is much less resource-intensive, requires significant engineering effort (Gerganov, 2023). Accelerating big Transformers remains an important open research question.
In the literature, Mixture of Experts (MoE)-based feedforward multi-layer perceptron (MLP) layers (Shazeer et al., 2017; Jacobs et al., 1991; Ivakhnenko & Lapa, 1965) have been popular methods to scale up Transformers to a large number of parameters (Lewis et al., 2021; Lepikhin et al., 2021; Fedus et al., 2022; Clark et al., 2022; Chi et al., 2022). However, in these works, the parameter efficiency of MoEs has not been studied; MoE models have been typically compared to dense baselines with the same number of FLOPs but with much less parameters. In fact, it has remained a common belief that they are not competitive against their dense counterparts with the same number of parameters. More recently, this parameter-efficiency perspective of MoEs has been studied by Csordás et al. (2023). They have proposed a novel MoE model, called $\sigma$-MoE, which not only speeds up Transformer training through reduced resource requirements, but also achieves performance comparable to or even superior to its parameter-matched dense counterparts; effectively demonstrating the promise of MoEs to scaling up neural networks.
Such MoE research often focuses on accelerating MLPs. However, attention layers (Schmidhuber, 1991; Bahdanau et al., 2015) also account for a considerable amount of compute and memory usage in Transformers, especially for long context sizes. Existing methods, such as linear attention (Schmidhuber, 1991; Katharopoulos et al., 2020; Choromanski et al., 2021; Peng et al., 2021; Schlag et al., 2021), typically fail to achieve significant speedups in practice (Dao et al., 2022) and often underperform its quadratic counterpart. As an alternative, MoE-based approaches for attention have been also proposed (Zhang et al., 2022; Peng et al., 2020), but in practice, they only achieve a modest reduction in computing and memory requirements, and typically require a lot of engineering tricks for successful training. Generally, MoE-based attention remains underexplored.
Here we present a novel MoE-based attention mechanism, SwitchHead, which reduces the number of attention matrices that need to be computed and stored. Our method is based on the $\sigma$-MoE by Csordás et al. (2023) and does not require regularization or extra tricks for stable training. We evaluate our method on a variety of language modeling datasets and two model sizes. We demonstrate that SwitchHead can achieve performance comparable to parameter-matched baselines, with a fraction of the compute and memory budget. Furthermore, we introduce the “SwitchAll” model, a fully MoE-based Transformer model, that combines a $\sigma$-MoE-based MLP layer with our SwitchHead attention, often outperforming dense baselines with the same parameter budgets. Lastly, we analyze the attention maps of our SwitchHead, and find that the maximum of attention maps taken over all heads are qualitatively similar to the dense baselines, indicating a significant reduction in redundancy without a loss of expressivity. Additionally, expert selections are often interpretable.
2 Method
Figure 1: Schematic representation of SwitchHead. It consists of few independent heads, each with multiple experts for value and output projections. Each head has a single attention matrix.
2.1 Background
The standard multi-head self-attention (MHA) layer (Vaswani et al., 2017) consists of four major steps: (1) computing key (K), query (Q), and value (V) projections, (2) computing the attention matrix, (3) using the attention matrix to project the values, and (4) mapping the projected values to the output. Let $h$, $T$, $d_{\text{model}}$, $d_{\text{head}}$ denote positive integers. Let ${\bm{x}}\in\mathbb{R}^{T\times d_{\text{model}}}$ denote an input to the MHA layer, $T$ be the sequence length, and $d_{\text{model}}$ denote the size of the hidden representations of the model. ${\bm{W}}_{\{K,V,Q\}}^{h}\in\mathbb{R}^{{d_{\text{model}}}\times{d_{\text{head}}}}$ are the projection matrices for head $h$. Then ${\bm{K}}^{h}={\bm{x}}{\bm{W}}_{K}^{h}$, ${\bm{Q}}^{h}={\bm{x}}{\bm{W}}_{Q}^{h}$, and ${\bm{V}}^{h}={\bm{x}}{\bm{W}}_{V}^{h}$ (thus ${\bm{K}}^{h},{\bm{Q}}^{h},{\bm{V}}^{h}\in\mathbb{R}^{T\times{d_{\text{head}}}}$) are the keys, queries, and values, respectively. The attention matrix for the head $h$, ${\bm{A}}^{h}\in\mathbb{R}^{T\times T}$, and the output ${\bm{y}}\in\mathbb{R}^{T\times{d_{\text{model}}}}$ are calculated as follows:
${\bm{A}}^{h}$ $=\mathrm{softmax}\left(\frac{1}{\sqrt{d_{\text{model}}}}{{\bm{Q}}^{h}{\bm{K}}^{h\intercal}}\right)$ (1)
${\bm{y}}$ $={\bm{W}}_{O}({\bm{A}}^{0}{\bm{V}}^{0}|{\bm{A}}^{1}{\bm{V}}^{1}|...|{\bm{A}}^{H}{\bm{V}}^{H})$ (2)
where |||| denotes concatenation in the last dimension, the $\mathrm{softmax}(\cdot)$ is also over the last dimension, and ${\bm{W}}_{O}\in\mathbb{R}^{d_{\text{model}}\times Hd_{\text{head}}}$. However, an alternative formulation reflects the role of ${\bm{W}}_{O}$ better. Let us divide ${\bm{W}}_{O}$ along the second dimension into submatrices for each head, ${\bm{W}}_{O}^{h}\in\mathbb{R}^{{d_{\text{model}}}\times d_{\text{head}}}$, such that ${\bm{W}}_{O}={\bm{W}}_{O}^{0}|{\bm{W}}_{O}^{1}|...|{\bm{W}}_{O}^{H}$. In this case, the output can be equivalently written as:
${\bm{y}}$ $=\sum_{h}{\bm{W}}_{O}^{h}{\bm{A}}^{h}{\bm{V}}^{h}$ (3)
From this, it can be seen that all computations are local to the heads. Computing the attention matrix $A^{h}$ and the readout $A^{h}V^{h}$ requires compute in order of $O(Hd_{\text{head}}T^{2})$ MACs (multiplication-accumulation operation). During training, it requires the storage of $O(HT^{2})$ for the attention matrices and $O(HTd_{\text{head}})$ numbers for storing the sub-results of the projections. Given a sufficiently long sequence, computing the attention matrix and projecting the values will dominate the compute requirements due to the quadratic dependence on the sequence length $T$.
2.2 From Dense to SwitchHead
Our goal is to obtain resource reductions while maintaining the fundamental properties of attention and retaining a fully expressive attention matrix. In fact, there is still room for improvement: modern LLMs use tens of heads (Brown et al., 2020; Touvron et al., 2023). Are so many of them all necessary? As we show later in Sec. 3, indeed, naively reducing the number of heads (while keeping the same number of parameters by increasing the head dimension) results in performance loss. Explaining the reason for the need for many heads is beyond the scope of this paper. Nevertheless, here are some hypotheses: (1) they provide multiple inputs for the operations that the network performs in each step, (2) they are specialized and provide inputs only for specific operations. In this case, each operation would use a different subset of heads. (3) They may also provide alternatives with different initializations, some being more successful than others, thus enabling better learning. Among these, some (2) and (3) offer an opportunity for resource savings: if not all heads are needed at the same time, it might be possible to switch between them. The simplest method of doing so is to produce a gating signal using a linear projection ${\bm{W}}_{S}\in\mathbb{R}^{d_{\text{model}}\times H}$, and use the ones with the highest activation, by replacing Eq. 3 with Eq. 6:
${\bm{s}}$ $=\sigma\left({\bm{x}}{\bm{W}}_{S}\right)$ (4)
$\mathcal{E}$ $=\operatorname*{arg\,topk}({\bm{s}},K),\mathcal{E}\subset\{1,...,H\}$ (5)
${\bm{y}}[t,c]$ $=\sum_{h\in\mathcal{E}}{\bm{s}}[t,h]({\bm{W}}_{O}^{h}{\bm{A}}^{h}{\bm{V}}^{h})[t,c]$ (6)
where ${\bm{y}}[t,c]$ denotes indexing the specific element of the matrix, specifically denoting timestep $t$ and channel $c$. Following Csordás et al. (2023), we use a non-competitive selection function. Intuitively, this corresponds to choosing a subset of attention heads for each output position. Our preliminary experiments confirmed that this method is indeed feasible for language modeling on WikiText-103. However, it is difficult to achieve acceleration and memory savings with this method. To see why, notice that the entries of the attention matrix ${\bm{A}}^{h}$ depend on pairs of inputs in different positions, but the choice is made only based on the output position. Thus, in the worst case, all possible projections have to be computed on the "source side" for the keys and values, which we would like to avoid.
An alternative approach, which we propose here, is to perform the conditional computation on the projections, independently for the source side ($K$ and $V$) and the destination side ($Q$ and output). This avoids conditional computation that involves the attention matrix itself. The obvious way to make the projections conditional is to use Mixture of Experts (MoEs). In this case, the concepts of "heads" are not well defined anymore. Therefore, we define a head to be a specific, computed attention matrix. For each head $h$, we define a list of $E$ experts. Then, the projection matrices become ${\bm{W}}_{K}^{h,e}$, ${\bm{W}}_{Q}^{h,e}$, ${\bm{W}}_{V}^{h,e}$ and ${\bm{W}}_{O}^{h,e}$, where $h$ denotes the head index and $e$ the specific expert. Then we compute the source-side expert selection as following:
${\bm{s}}_{S}^{h}$ $=\sigma({\bm{x}}{\bm{W}}_{S}^{h})$ (7)
$\mathcal{E}_{S}^{h}$ $=\operatorname*{arg\,topk}({\bm{s}}_{S}^{h},K),\mathcal{E}_{S}^{h}\subset\{1,...,E\}$ (8)
We compute the destination-side experts similarly: ${\bm{s}}_{D}^{h}=\sigma({\bm{x}}{\bm{W}}_{D}^{h})$, $\mathcal{E}_{D}^{h}=\operatorname*{arg\,topk}({\bm{s}}_{D}^{h},K),\mathcal{E}_{S}^{h}\subset\{1,...,E\}$. Then, the value projection ${\bm{V}}^{h}$ is computed as a weighted sum of the selected experts:
${\bm{V}}^{h}$ $=\sum_{e\in\mathcal{E}_{S}^{h}}{\bm{s}}_{S}^{h}[e]{\bm{x}}{\bm{W}}_{V}^{h,e}$ (9)
The key and query projections are computed similarly: ${\bm{K}}^{h}=\sum_{e\in\mathcal{E}_{S}^{h}}{\bm{s}}_{S}^{h}[e]{\bm{x}}{\bm{W}}_{K}^{h,e}$, and ${\bm{Q}}^{h}=\sum_{e\in\mathcal{E}_{D}^{h}}{\bm{s}}_{D}^{h}[e]{\bm{x}}{\bm{W}}_{Q}^{h,e}$. The output projection also becomes an MoE:
${\bm{y}}$ $=\sum_{h=0}^{H-1}\sum_{e\in\mathcal{E}_{D}^{h}}{\bm{W}}_{O}^{h,e}{\bm{A}}^{h}{\bm{V}}^{h}$ (10)
As we’ll show, it is not necessary to make all projections MoEs. In Section 3.1 we show that keeping a single copy of the projections $Q$ and $K$ and reusing them for all experts is beneficial. We call this method SwitchHead. If this method can reduce the number of heads $H$ by having more experts, $E$, then it provides an easy way to reduce the resource requirements of MHA. Note that our method does not depend on the specific implementation of the attention, allowing easy experimentation and research. A schematic representation is shown in Fig. 1.
Unlike standard MoE methods, we found that no regularization is necessary to achieve good performance with our method.
2.3 Resource Usage of Different Methods
In this section, we discuss the compute and memory usage of different attention variants. We will define the compute in terms of the number of multiply-accumulate operations (MACs, also used by Zhang et al. (2022)), which is arguably better defined than FLOPs (e.g., does one step of the matrix multiplication count as 1 FLOP or 2? Do we include the softmax?). All calculations will be presented for a single attention layer for a single sequence, and they are presented this way in all our tables. Both the memory and compute requirements scale linearly with both the batch size and the number of layers.
Consider a sequence of inputs of length $T$, with representation size $d_{\text{model}}$. Let $d_{\text{head}}$ be the width of the $K$, $Q$, and $V$ projections used for the attention layer. For Transformer XL-style attention, let the size of the context be $CT$, where $C-1$ is the number of past chunks included in the context of the current attention step. We can divide the computation into two major parts: calculating the projections, which do not involve the attention map, and calculating the attention map and projecting the sequence of values using it.
First, consider the case of the standard Transformer XL (Dai et al., 2019). Here, from the input ${\bm{x}}\in\mathbb{R}^{T\times d_{\text{model}}}$, we calculate the ${\bm{K}}^{h},{\bm{Q}}^{h},{\bm{V}}^{h}\in\mathbb{R}^{T\times{d_{\text{head}}}}$ using projection matrices of shape $\mathbb{R}^{d_{\text{model}}\times d_{\text{head}}}$. The output after the attention is projected in a similar manner (Eq. 3). Thus, the projections take a total of $4Td_{\text{model}}d_{\text{head}}$ MACs per head. For backpropagation, we have to store all the intermediate results. This takes $Td_{\text{head}}$ numbers of ${\bm{K}}^{h}$, ${\bm{Q}}^{h}$ and ${\bm{V}}^{h}$. Also, the projected values should be stored. They have an identical shape, therefore, the total memory used by projections is $4Td_{\text{head}}$ numbers per head. Now consider the resource usage related to the attention matrix. It involves calculating the product of ${\bm{Q}}^{h}{\bm{K}}^{h\intercal}$, which takes ${d_{\text{head}}}CT^{2}$ MACs (multiplication by $C$ is needed because the shape of ${\bm{K}}^{h}$ and ${\bm{V}}^{h}$ for Transformer XL is $CT\times d_{\text{head}}$). The projection of the values with the attention matrix ${\bm{A}}^{h}{\bm{V}}^{h}$ is similar. For the memory usage, the attention needs $CT^{2}$ numbers, but it needs to be stored both before and after the activation function. In addition, calculating the projection of the position encodings is necessary. This depends on the implementation, but in our case, it involves a matrix multiplication, and the total amount of computation is $2{d_{\text{head}}}{d_{\text{model}}}TC$, and it needs $2{d_{\text{head}}}TC$ numbers of storage. Thus the resource requirements are:
$N_{\text{MAC}}^{\text{XL}}$ $=H\left(4Td_{\text{head}}d_{\text{model}}+2CT^{2}d_{\text{head}}+2CTd_{\text{head}}d_{\text{model}}\right)$ (11)
$N_{\text{mem}}^{\text{XL}}$ $=H\left(4Td_{\text{head}}+2CT^{2}+2CTd_{\text{head}}\right)$ (12)
The resource usage of SwitchHead is different. First, the number of heads $H$ is significantly reduced, but $d_{\text{head}}$ is typically larger. Additionally, there are $K$ experts active at the same time. Here, we only consider the case where the value and outputs are experts, but ${\bm{Q}}^{h}$ and ${\bm{K}}^{h}$ are not (this version performs the best; see Sec. 3.1). Then, we have two projections that are identical with that of Transformer XL, and two MoE-based projections. These use $TKd_{\text{model}}d_{\text{head}}$ MACs to calculate the projection and another $TKd_{\text{head}}$ to calculate their weighted average. With a smart kernel implementation, memory usage is not affected by $K$, thus the formula remains the same as Eq. 12 (note, however, that $H$ and $d_{\text{head}}$ are very different in practice). The compute requirement can be calculated as:
$N_{\text{MAC}}^{\text{{SwitchHead}}}$ $=H\left(2Td_{\text{head}}d_{\text{model}}+2TKd_{\text{head}}(d_{\text{model}}+1)+2CT^{2}d_{\text{head}}+2CTd_{\text{head}}d_{\text{model}}\right)$ (13)
Additionally, the expert selection logic needs minimal additional resources, which can be ignored. Note that the comparison between the MACs of the standard (Eq. 11) and SwitchHead (Eq. 13) depends on the exact values of the hyper-parameters. However, as we’ll see in Sec. 3, in our typical configurations, SwitchHead provides good predictive performance with significantly lower $H$ compared to the standard Transformer, resulting in reduced resource usage in the end.
3 Experiments
Following Csordás et al. (2023) we conduct our experiments in a parameter-matched setting which better reflects the expressivity of language models (than the FLOPS-matched setting often used to evaluate MoEs). Without this constraint, with MoEs it is often possible to compensate for a weaker method by adding more experts. We use and adopt the CUDA kernel of Csordás et al. (2023) for our purposes. To match the number of parameters of different models, we follow a systematic procedure. First, we measure the parameter count of the dense Transformer, which serves as our target. Then, for each method, we set the total number of experts (including between heads, $HE$ for SwitchHead) to the same as the original number of heads. We increase the head projection size $d_{\text{head}}$ to the maximum that keeps the parameter count below our target. Because our CUDA kernel supports only $d_{\text{head}}$ with multiples of 4, this often remains below the parameter count of the baseline. For further compensation, we slightly increase $d_{\text{ff}}$ until we achieve a match that differs from our target with no more than 100k parameters but never exceeds it. We do not claim that this parameter-matching method is optimal, but we aim to have a consistent algorithm that does not require tuning, which is prohibitively expensive and would have to be done for each model separately. Detailed hyperparameters of all our models can be found in Sec. A.2 in the Appendix.
For all datasets except the character-level Enwik8 (Hutter, 2006), we use sub-word units (Sennrich et al., 2016; Schuster & Nakajima, 2012) obtained with a SentencePiece tokenizer (Kudo & Richardson, 2018) with a vocabulary size of 8k tokens. Unless otherwise noted, all models, including ours, are Transformer XL (Dai et al., 2019), with the context size being twice the size of the active/current chunk.
All models are trained for 100k batches. Some of the datasets we consider (C4 (Raffel et al., 2020), and peS2o (Soldaini & Lo, 2023)) are much larger. In this case, we train on the first $10^{5}*T*N_{\text{batch}}$ tokens of the dataset.
3.1 Which Projections Require an MoE?
As discussed in Sec. 2.2, each linear projection (K, V, Q, O) can potentially be replaced by an MoE. Here we first check which projection benefits from such a replacement. As we target the parameter-matched setting, having experts where they are not necessary can have a negative effect. Since they use a significant part of the parameter budget, they can reduce the number of parameters available for the more useful parts of the model. Thus, we did a search over all possible combinations of expert versus fixed projections with two active heads and compared them to the parameter-matched baseline on Wikitext 103. Our models have 47M parameters. We also include a parameter-matched baseline with two heads, which serves as a lower bound for the performance. The results are shown in Tab. 1. It can be seen that the output projection is necessary to match the performance of the baseline. Having key and query experts seems to be unnecessary. In fact, without the output and value experts, they even underperform the dense baseline with $H=2$ heads. The best-performing model is the one with experts for both value and output projections. We use this model variant for all the other experiments in this paper.
Table 1: Performance of SwitchHead with $E=5$ experts and $H=2$ heads. Different projections are either experts or fixed for the given head. Parameter-matched baseline with $H=10$ and $H=2$ are shown. Models sorted by perplexity. 47M parameters models on Wikitext 103.
Model $n_{\text{heads}}$ V expert K expert Q expert O expert Perplexity
SwitchHead 2 Y N N Y 12.27
SwitchHead 2 N N N Y 12.30
Transformer XL 10 - - - - 12.31
SwitchHead 2 N Y N Y 12.36
SwitchHead 2 Y Y N Y 12.37
SwitchHead 2 Y N Y Y 12.42
SwitchHead 2 Y N N N 12.45
SwitchHead 2 N N Y Y 12.45
SwitchHead 2 Y N Y N 12.51
SwitchHead 2 Y Y Y Y 12.57
SwitchHead 2 N Y Y Y 12.59
SwitchHead 2 Y Y Y N 12.61
SwitchHead 2 Y Y N N 12.69
Transformer XL 2 - - - - 12.74
SwitchHead 2 N N Y N 12.75
SwitchHead 2 N Y N N 12.79
SwitchHead 2 N Y Y N 12.90
3.2 Comparing with MoA
The method most related to ours is the so-called Mixture of Attention Heads, or MoA (Zhang et al., 2022). They use a selection mechanism to choose active attention heads from a set of experts. However, they have a single set of $K$ and $V$ projections shared between experts; thus, acceleration is possible. However, in the original paper, the authors use a high number of selected heads (8-16) which seems necessary to achieve good performance. Thus, the resource reductions are moderate. Moreover, MoA uses three different regularizers, which have to be tuned independently.
We compare our method with our reimplementation of MoA with a different number of selected heads. Given the complexity of tuning its regularization coefficients, we take them directly from Zhang et al. (2022). For a fair comparison, we also integrated the non-competitive selection mechanism of Csordás et al. (2023) into MoA. The results are shown in Table 2. Similarly to our method, we found that with non-competitive selection, no regularization is required, and the predictive performance usually is superior to the original formulation. However, it still underperforms our method given a similar computation and memory budget.
Table 2: Performance of SwitchHead compared to different MoA variants. "sel. activation" denotes the activation function used for selecting the experts. the MoA can outperform the baseline, but only at a price of using significantly more computing and memory. Also, SwitchHead outperforms the baseline dense Transformer. Results are on Wikitext 103.
Model sel. activation $n_{\text{heads}}$params Perplexity MACs Mem (floats)
MoA sigmoid 8 47M 12.13 390.2M 2.6M
MoA sigmoid 6 47M 12.16 306.8M 1.9M
SwitchHead sigmoid 2 47M 12.27 170.4M 0.8M
Transformer XL - 10 47M 12.31 453.4M 3.5M
MoA sigmoid 4 47M 12.39 223.5M 1.3M
MoA softmax 4 47M 12.60 223.5M 1.3M
MoA softmax 6 47M 12.64 306.8M 1.9M
MoA sigmoid 2 47M 12.65 140.1M 0.7M
MoA softmax 8 47M 12.77 390.2M 2.6M
MoA softmax 2 47M 12.84 140.1M 0.7M
MoA softmax 8 262M 9.50 2.9G 9.9M
SwitchHead sigmoid 2 262M 9.55 2.0G 2.9M
MoA sigmoid 8 262M 9.56 2.9G 9.9M
MoA sigmoid 12 262M 9.58 4.1G 14.7M
Transformer XL - 16 262M 9.66 5.4G 21.0M
MoA softmax 12 262M 9.68 4.1G 14.7M
MoA softmax 4 262M 9.69 1.7G 5.1M
MoA sigmoid 4 262M 9.77 1.7G 5.1M
MoA softmax 2 262M 9.87 1.1G 2.7M
MoA sigmoid 2 262M 10.02 1.1G 2.7M
3.3 Performance on Different Datasets
We test our methods on a diverse set of language modeling datasets, including C4 (Raffel et al., 2020), Enwik8 (Hutter, 2006), peS2o (Soldaini & Lo, 2023), at two different scales: a 47M and a 262M parameters. The results are shown in Tab. 3. We compare our models to two baselines: one with the same number of heads as the total number of experts ($H\cdot E$) of the SwitchHead models, and the other has the same number of heads as the number of active attention matrices ($H$) as our models. Our models always closely match the performance of the full, many-head baseline with the fraction of memory and compute requirements. Importantly, our method also achieves a wall-clock speedup, enough to accelerate the entire training pipeline by a factor of around 1.5 (see Appendix A.4 for more details). This confirms the competitiveness of our method.
Table 3: Performance of SwitchHead compared to baselines on different datasets with different model sizes. It can be seen that the predictive performance of our SwitchHead model is comparable to the baselines, and is always better than the baseline with an equal number of heads. Perplexity is shown for Wikitext 103, C4 and peS2o datasets, and bits/character (bpc) for Enwik8.
Model Dataset $n_{\text{heads}}$params ppl/bpc MACs Mem (floats)
SwitchHead C4 2 47M 22.49 202.5M 0.8M
Transformer XL C4 10 47M 22.63 453.4M 3.5M
Transformer XL C4 2 47M 23.71 453.4M 1.4M
SwitchHead C4 4 262M 16.38 2.4G 5.6M
Transformer XL C4 16 262M 16.58 5.4G 21.0M
Transformer XL C4 4 262M 17.09 5.4G 8.4M
SwitchHead Wikitext 103 2 47M 12.31 170.4M 0.8M
Transformer XL Wikitext 103 10 47M 12.32 453.4M 3.5M
Transformer XL Wikitext 103 2 47M 12.73 453.4M 1.4M
SwitchHead Wikitext 103 2 262M 9.77 2.0G 2.9M
Transformer XL Wikitext 103 16 262M 9.80 5.4G 21.0M
Transformer XL Wikitext 103 2 262M 10.09 5.4G 6.3M
Transformer XL peS2o 10 47M 12.83 453.4M 3.5M
SwitchHead peS2o 2 47M 12.84 202.5M 0.8M
Transformer XL peS2o 2 47M 13.37 453.4M 1.4M
Transformer XL peS2o 16 262M 9.78 5.4G 21.0M
SwitchHead peS2o 4 262M 9.86 2.4G 5.6M
Transformer XL peS2o 4 262M 10.11 5.4G 8.4M
SwitchHead Enwik8 2 41M 1.10 709.3M 2.8M
Transformer XL Enwik8 8 41M 1.10 1.6G 10.5M
Transformer XL Enwik8 2 41M 1.13 1.6G 4.2M
3.3.1 SwitchAll
The goal of achieving more resource-efficient Transformers includes reducing the resource requirements of both the MLP and the attention layers. Csordás et al. (2023) proposed a parameter-efficient MoE method to accelerate the MLP layers. However, it remains unclear whether it can be efficiently combined with our SwitchHead, or can have some negative interaction effect if combined in a "SwitchAll", where every layer is MoE-based.
In order to verify this, we take the architecture proposed by Csordás et al. (2023) without any hyperparameter change and replace the attention layer with SwitchHead. The hyperparameters for the attention are directly taken from the experiments shown in Tab. 3. The results are shown in Tab. 4. The combined, fully-MoE model often outperforms the dense baselines for each dataset and model size considered, except in the case of the 259M parameter model on the C4 dataset.
Table 4: Performance of SwitchAll (SwitchHead + $\sigma$-MoE (Csordás et al., 2023)) on different datasets and model sizes. Our SwitchAll model is close or better compared to the baselines.
Model Dataset $n_{\text{heads}}$params ppl/bpc MACs Mem (floats)
SwitchAll Wikitext 103 2 47M 12.17 170.4M 0.8M
Transformer XL Wikitext 103 10 47M 12.32 453.4M 3.5M
Transformer XL Wikitext 103 16 262M 9.80 5.4G 21.0M
SwitchAll Wikitext 103 4 259M 9.81 2.4G 5.6M
SwitchAll C4 2 47M 22.09 202.5M 0.8M
Transformer XL C4 10 47M 22.63 453.4M 3.5M
SwitchAll C4 4 259M 16.45 2.4G 5.6M
Transformer XL C4 16 262M 16.58 5.4G 21.0M
SwitchAll peS2o 2 47M 12.56 202.5M 0.8M
Transformer XL peS2o 10 47M 12.83 453.4M 3.5M
Transformer XL peS2o 16 262M 9.78 5.4G 21.0M
SwitchAll peS2o 4 259M 9.86 2.4G 5.6M
4 RoPE Positional Encodings
All of our experiments so far have used a Transformer XL model. Thus, it remains unclear whether SwitchHead is specific to this model or can be also used with other attention methods. As an alternative, we consider RoPE positional encodings Su et al. (2021) without the XL cache (thus, the attention matrices are square). We test these models on Wikitext 103. The results are shown in Tab. 5. Our method also performs well in this case.
Table 5: Performance of SwitchHead compared to dense baseline on Wikitext 103, using RoPE positional encoding instead of Transformer XL.
Model Dataset $n_{\text{heads}}$params ppl/bpc MACs Mem (floats)
SwitchHead (RoPE) Wikitext 103 2 45M 12.75 285.6M 1.3M
Transformer (RoPE) Wikitext 103 10 45M 12.78 560.9M 6.1M
Transformer (RoPE) Wikitext 103 2 45M 12.96 560.9M 1.9M
SwitchHead (RoPE) Wikitext 103 4 243M 10.00 4.2G 18.4M
Transformer (RoPE) Wikitext 103 16 244M 10.17 6.4G 37.7M
Transformer (RoPE) Wikitext 103 2 244M 10.26 6.4G 8.4M
5 Analysis
In order to see how the network uses the attention heads, we trained a small, 6-layer, 8-head Transformer on ListOps (Nangia & Bowman, 2018; Csordás et al., 2022). The reason for this choice is that small, algorithmic tasks tend to be more interpretable compared to language models. We also train a parameter-matched, 2-head SwitchHead model. Both models achieve around 95% accuracy on a held-out IID validation set, in contrast to the dense 2-head model, which saturates around 80%. Note that ListOps is a classification task and does not use autoregressive masking.
Following Csordás et al. (2022), we visualize the maximum of attention heads for each layer, both for the standard Transformer (Fig. 2a) and SwitchHead (Fig. 2b). The attention maps are qualitatively similar. Note that the initialization and the learning dynamics are different for the two models, thus the overlap would not be perfect even with the same type of model. We show all the attention maps for both models in Fig. 4 and 3 in the Appendix.
In addition, we visualize individual attention heads for the SwitchHead model. An example is shown in Fig. 2c. In addition to the attention map, we show the weight of the selected experts for both the value and output projection (denoted by V and O, respectively, on the sides of the attention map). Often it is possible to interpret the selection weights: here, the output experts specialize according to different operations, while the input ones distinguish numbers and closed parentheses. The attention map itself appears to distribute information about contiguous chunks of numbers. Similar plots for all heads are shown in Fig. 5 in the Appendix.
(a) Transformer, Layer 3
(b) SwitchHead Layer 3
(c) SwitchHead, Layer 3, Head 1
Figure 2: An attention map of the (a) standard Transformer and (b) SwitchHead. The maximum of all heads in the given layer are shown. (c) A head of SwitchHead. On the left side of the attention plot, the selection weights of the output projection expert are shown. Similarly, at the bottom, the selection weights of the value experts are visible. In the selection maps, dark blue always corresponds to 1, while white is 0. The scale shown on the right is only for the attention.
The attention maps of the language models are difficult to interpret. However, we visualized the attention maps of the 47M parameter Transformer XL and the SwitchHead model from Tab. 3. We found them to be qualitatively similar. We also identified induction heads (Olsson et al., 2022) in both models, some examples shown for SwitchHead in Fig. 6a and for Transformer in Fig. 6b in the appendix. Other typical vertical line-lined attention patterns are shown in Fig. 6c and 6d.
6 Related Work
The method most closely related to ours is MoA (Zhang et al., 2022), which introduces a MoE style attention. It defines each attention head as an expert but shares the key and value projections between them. Unlike in our case, each of the selected experts requires a separate attention matrix, which significantly increases its memory usage. Due to the use of a competitive softmax-based activation function in the selection network, it requires complex regularization to prevent expert collapse. In the original formulation, the number of active heads is high. We also confirmed in our experiments that MoA needs many attention heads to match the performance of the dense baseline (see Sec. 3.2), and it is only possible to do so with a significantly higher resource budget than our method.
Nguyen et al. (2022) analyze the attention matrices, and they conclude that they are usually low rank. Motivated by this, the authors construct a few (e.g., 2) "global attention matrices", and they compute each local matrix for specific heads by a weighted average of those. However, they average the logits, not the final matrix, so each individual head-specific matrix has to be computed. This means that in the best case, they can only save half of the computation associated with the attention matrix because the readout (Eq. 3) is still needed. For the same reason, memory savings are also low. The authors also use sampling of the attention matrices.
Peng et al. (2020) proposes to reweight the contribution of each head by a gating function. However, they only reduce the number of total attention heads by one, presumably to compensate for the parameters used by the selection logic. Their goal was not to reduce resource usage but to have better predictive performance, which they achieve. They use a softmax-based competitive selection mechanism. To avoid collapse, the gating function is trained only in some steps.
Csordás et al. (2023) introduce the non-competitive $\sigma$-MoE method that we also use for our attention mechanism. However, the authors focus on accelerating the MLPs and not the attention. More broadly, Shazeer et al. (2017) introduces sparsely-gated mixture of experts in LSTM (Hochreiter & Schmidhuber, 1997) networks. Fedus et al. (2021) introduces Mixture of Experts in Transformers. Lepikhin et al. (2021) trains a MoE-based LLM, and Clark et al. (2022) analyzes the scaling laws of MoE models. Lewis et al. (2021) introduces an alternative method for preventing collapse.
Dao et al. (2022) provides a hardware-aware CUDA implementation of the entire attention layer, which avoids storing the attention matrix. By saving memory bandwidth in this way, they achieve a significant wall clock time speedup, despite that the attention matrix should be recomputed in the backward pass. This is orthogonal to our method and they can be combined for further acceleration.
7 Conclusion
On a wide range of language modeling datasets with different model sizes, our novel Mixture-of-Experts-based attention method called SwitchHead achieves performance on par with parameter-matched dense counterparts, but with only a fraction of the computational cost and memory usage. SwitchHead drastically reduces the number of attention matrices that have to be computed, by using MoE for the value and output projections. Our method is stable and does not need additional regularization to prevent degenerate solutions (a well-known practical issue in many existing MoE models). Our method can also be successfully combined with MoE MLP layers, to obtain "SwitchAll" where every layer of the Transformer is MoE-based, achieving a huge reduction in resource requirements.
Acknowledgements
This research was partially funded by ERC Advanced grant no: 742870, project AlgoRNN, and by Swiss National Science Foundation grant no: 200021_192356, project NEUSYM. We are thankful for hardware donations from NVIDIA and IBM. The resources used for this work were partially provided by Swiss National Supercomputing Centre (CSCS) projects d123 and s1205.
Appendix A Appendix
A.1 Resource Requirements of MoA
The resource requirements of MoA (Peng et al., 2020) are very similar to those of Transformer XL (see Sec. 2.3 for more details), except that it uses a single shared ${\bm{K}}$ and ${\bm{V}}$ for each head.
$N_{\text{MAC}}^{\text{MoA}}$ $=(2H+2)Td_{\text{head}}d_{\text{model}}+2HCT^{2}d_{\text{head}}+2CTd_{\text{head}}d_{\text{model}}$ (14)
$N_{\text{mem}}^{\text{MoA}}$ $=(2H+2)Td_{\text{head}}+2HCT^{2}+2CTd_{\text{head}}$ (15)
A.2 Hyperparameters
We train all our models with Adam optimizer (Kingma & Ba, 2015), with a batch size of 64, a learning rate of 0.00025, and gradient clipping with a maximum norm of $\kappa$. Large models ($>200K$ parameters) use a learning rate warm-up of 4k steps. All models, except the SwitchAll model, use a dropout on the MLP layers, $0.1$ for the small models and $0.2$ for the large ones. Detailed hyperparameters are shown in the Tab. 6. $\sigma$-MoE related hyperparameters for the SwitchAll models are identical to those of Csordás et al. (2023). For Transformer XL models, we always use a single additional chunk of context, both in training and validation time. $d_{\text{head}}$ and $d_{\text{ff}}$ are derived in a systematic way, see Sec. 3 for more details.
Table 6: Hyperparameters used for our models.
Model Dataset $n_{\text{heads}}$params $d_{\text{head}}$ $d_{\text{ff}}$ E K T $n_{\text{layers}}$ $\kappa$
SwitchHead C4 2 47M 76 2080 5 3 256 16 0.1
Transformer XL C4 10 47M 41 2053 - - 256 16 0.1
Transformer XL C4 2 47M 205 2053 - - 256 16 0.1
SwitchHead C4 4 262M 112 4188 4 2 512 18 0.25
Transformer XL C4 16 262M 64 4110 - - 512 18 0.25
Transformer XL C4 4 262M 256 4110 - - 512 18 0.25
SwitchHead Wikitext 103 2 47M 76 2080 5 2 256 16 0.1
Transformer XL Wikitext 103 10 47M 41 2053 - - 256 16 0.1
Transformer XL Wikitext 103 2 47M 205 2053 - - 256 16 0.1
SwitchHead Wikitext 103 2 262M 132 4147 8 4 512 18 0.25
Transformer XL Wikitext 103 16 262M 64 4110 - - 512 18 0.25
Transformer XL Wikitext 103 2 262M 512 4110 - - 512 18 0.25
SwitchHead peS2o 2 47M 76 2080 5 3 256 16 0.1
Transformer XL peS2o 10 47M 41 2053 - - 256 16 0.1
Transformer XL peS2o 2 47M 205 2053 - - 256 16 0.1
SwitchHead peS2o 4 262M 112 4188 4 2 512 18 0.25
Transformer XL peS2o 16 262M 64 4110 - - 512 18 0.25
Transformer XL peS2o 4 262M 256 4110 - - 512 18 0.25
SwitchHead Enwik8 2 41M 112 2088 4 2 512 12 0.25
Transformer XL Enwik8 8 41M 64 2053 - - 512 12 0.25
Transformer XL Enwik8 2 41M 256 2053 - - 512 12 0.25
SwitchHead (RoPE) Wikitext 103 2 45M 64 2092 5 3 512 16 0.1
Transformer (RoPE) Wikitext 103 10 45M 41 2053 - - 512 16 0.1
SwitchHead (RoPE) Wikitext 103 4 243M 100 4136 4 2 1024 18 0.25
Transformer (RoPE) Wikitext 103 16 244M 64 4110 - - 1024 18 0.25
SwitchAll Wikitext 103 2 47M 76 1648 5 2 256 16 0.25
SwitchAll Wikitext 103 4 259M 112 4096 4 2 512 18 0.25
SwitchAll C4 2 47M 76 1648 5 3 256 16 0.25
SwitchAll C4 4 259M 112 4096 4 2 512 18 0.25
SwitchAll peS2o 2 47M 76 1648 5 3 256 16 0.25
SwitchAll peS2o 4 259M 112 4096 4 2 512 18 0.25
A.2.1 A Note on Choosing $n_{\text{heads}}$
Our preliminary experiments showed that a single head is usually not enough to match the performance of the baseline network, but two heads usually work well. Because of this, we always start by training a model with $n_{\text{heads}}=2$ and increase it to $n_{\text{heads}}=4$ if it does not match the performance of the baseline. We have not experimented with any other $n_{\text{heads}}$.
A.3 A Note on the Parameter Count of the SwitchAll
It can be seen in Tab. 4 that the parameter count of the SwitchAll models is often less than that of their dense counterparts. The reason is that we normally compensate for the final difference in the number of parameters by increasing $d_{\text{ff}}$ (see Sec. 3 for details of the parameter matching). However, that can only be done in a very coarse-grained way with $\sigma$-MoE: the size of all experts must be increased at once, and the CUDA kernel supports only sizes of multiple of 4. Therefore, increasing the size of the experts would add too many parameters and the model would outgrow the baseline. For this reason, we simply keep the hyperparameters for Csordás et al. (2023) and combine them with our SwitchHead configuration from Tab. 3.
A.4 Wall-Clock Time Estimation
In all of our tables, we report the number of multiply-accumulate (MAC) operations following Zhang et al. (2022). The reason for this is that the actual wall-clock time is highly implementation and hardware-dependent. Nevertheless, we measured the runtime and total memory usage of our entire training pipeline (including the feedforward layer) to demonstrate that our current (suboptimal) implementation is already capable of providing wall-clock-time acceleration. We show the results in Tab. 7. The measurements are taken on identical hardware with the same implementation (including for the attention core), the only difference being the MoE-based projections for the attention. It can be seen that for both scales, our method trains around 1.5 times faster, while using 61%-67% as much memory as the baseline. Note that these measurements also include the MLP layers, the optimizer, and the gradient synchronization in the case of multi-GPU training.
Table 7: Real-world resource usage of our method. The numbers shown below are for training time for the whole pipeline, including the feedforward layers. It can be seen that SwitchHead in the current implementation reduces both the runtime and the memory usage by a factor of 1.4-1.5.
Model Size ms/iteration Rel. iter. time RAM/GPU Rel. Mem. #GPUs GPU type
Trafo. XL 47M 473ms/iter 1.0 20.5G 1.0 1 RTX 3090
SwitchHead 342ms/iter 0.72 13.5G 0.65
Trafo. XL 262M 670ms/iter 1.0 20.5G 1.0 8 V100
SwitchHead 442ms/iter 0.65 12.5G 0.61
A.5 Visalizing all Attention Heads
As discussed in Sec. 5, we analyze the attention maps of SwitchHead and compare them with the dense models. We show all the attention maps of the models trained on ListOps in Fig. 3 and Fig. 3. We show individual heads of SwitchHead, including the expert selection scores in Fig. 5. Some selected attention maps of our 47M parameter models on Wikitext 103 are shown in Fig. 6.
(a) Layer 1
(b) Layer 2
(c) Layer 3
(d) Layer 4
(e) Layer 5
(f) Layer 6
Figure 3: The maximum of all attention maps for a SwitchHead model on ListOps.
(a) Layer 1
(b) Layer 2
(c) Layer 3
(d) Layer 4
(e) Layer 5
(f) Layer 6
Figure 4: The maximum of all attention maps for a standard Transformer model on ListOps.
(a) Layer 1, head 1
(b) Layer 1, head 2
(c) Layer 2, head 1
(d) Layer 2, head 2
(e) Layer 3, head 1
(f) Layer 3, head 2
(g) Layer 4, head 1
(h) Layer 4, head 2
(i) Layer 5, head 1
(j) Layer 5, head 2
(k) Layer 6, head 1
(l) Layer 6, head 2
Figure 5: Details for individual heads of the SwitchHead model on ListOps. On the left side of each attention plot, the selection of the output projection expert is shown. Similarly, at the bottom, the selection of the value projection selection is visible. In the selection maps, dark blue always corresponds to 1, while white is 0. The adaptive scale shown to the right of the attention map is for the map only.
(a) SwitchHead Layer 12. Induction head.
(b) Transformer XL Layer 10. Induction head.
(c) SwitchHead Layer 9. Stripe pattern.
(d) Transformer XL Layer 8. Stripe pattern.
Figure 6: Induction head copying the rare name "Homarus" in (a) SwitchHead and (b) Transformer XL baseline. The attention matrix is square because it is the first chunk of the sequence, without any extra context. Typical vertical line pattern in (c) SwitchHead and (b) Transformer XL baseline.
n8HAOVy+T+XLJfLS4ZhTiRJgqIoVBRFIoric47jPnmeB1mW/9rr9ZpSSn3Lsmir1fJZlqWlUonKsvwWwD8ymc/nXwVBeLjf7xEKhdBut9Hr9WgmkyGEkJwsy5eHG5vN5g0AKIoCAEgkEkin0wQAfN9/cXPdheu6P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">